# Deformable Cluster Manipulation via Whole-Arm Policy Learning

### v1

@@ Abstract
Videos are available @ https://youtu.be/hG-zYoesX5k
@@ I INTRODUCTION
Learning to manipulate deformable objects poses a formidable challenge in robotics [1][2]; however, operating with clusters of deformables, such as cable bundles, multi-branch tree canopies, and cloth piles, is significantly harder. Although imprecise physics knowledge [3] and sensor uncertainties are contributing factors, the principal challenge lies in inferring a compact, low-dimensional state abstraction and the temporal state evolution, based on noisy perception from a high-dimensional ground truth [4]. Furthermore, with deformable clusters, tiny differences in contact locations, material properties, and initial configuration among the individual cluster constituents (e.g. between thicker and thinner plant stems) give rise to non-linear resistance forces and divergent motion trajectories upon interaction. The resulting chaotic dynamics render model-based learning infeasible, while model-free solutions have high sample complexity, and therefore, the compactness and generation efficiency of the state representation gather significance [5].
@@
Novel deformable works manipulating cables [6], dough [7], and clothes [8] are all planning approaches; furthermore, they exclusively focus on end-effector-based control. In a cluster context, such grasp-based solutions tailored to individual deformable members become prohibitively long. On the contrary, a multi-link manipulation strategy substantially escalates the contact complexity, necessitating additional perception modalities apart from vision. However, typical auxiliary sensors for whole-arm contact detection are either expensive [9] or too slow [10] for low-latency applications.
Addressing these challenges, we propose a model-free, non-prehensile, multi-modal, reinforcement learning approach (Fig. 1) operating on the cluster as a whole with full-body contact, but without external tactile sensing. We use a distributional representation [4] of the cluster state generated by embedding point cloud samples into a Reproducing Kernel Hilbert Space (RKHS), without key-point intermediaries. Compared to traditional point cloud feature extractors, a distribution embedding is notably lightweight, enabling faster training and real-time reinforcement learning (RL) feedback. It eliminates the need for points to track identical cluster locations, is invariant to permutations, and can handle variable input sizes resulting from cluster members moving out of camera view between time steps. Further, the observations include a whole-arm contact classifier [11] output from proprioceptive measurements to flag contact. Finally, we address the high sample requirements of on-policy RL by leveraging the parallelism of physics simulators for policy training and transferring it zero-shot to a real-world setting with completely unseen geometry and dynamics.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d)
Figure 2: (a) Manual pruning of overhanging branches impeding power lines, from [12] (b) Network cables occluding autonomous server inspection (c) Our laboratory setting mimicking the power line clearance scenario (d) Simulation environment in Isaac Gym with a parametric L-system ternary structure.
This work focuses on a specific subclass of problems, aiming to de-occlude local regions of the deformable cluster; for instance, to clear a cylinder-shaped region among tree branches, by manipulating a few of them together. Illustrative applications of such a skill abstraction include:
@@
    Clearing Power Lines: Hazardous tree branches encroaching on overhead power infrastructure is a leading cause of forest fires and outages [13][14]. Pruning foliage near overhead lines (Fig. 2a). is dangerous for humans due to risks like falls, shocks, and tool accidents [12]; therefore, robotic assistive systems that can clear the branches away from the lines before pruning can enhance safety while reducing outages. In this instance, the cylindrical clearance region can be thought of as concentric to the physical power line.
@@
    Autonomous Inspection: Modern electrical and networking infrastructure, such as a data centre, is cluttered with deformable electrical and network cables obstructing inspection cameras (Fig. 2b). Autonomous monitoring robots must flexibly manipulate cable bunches to restore camera line-of-sight for back panel inspection and anomaly detection. Here, the clearance region may be a virtual cylinder with the inspection camera on one face and the target panel on the other.
  3. 3.
    Agricultural Exposure: Modern harvesting systems suffer from significantly lower success rates with partially (50%-75%) and fully occluded (5%) fruits [15]. Manipulating plant stems for exposing fruits or a diseased region to an external sensor requires a line-of-sight restoration through multi-branch clearance.
Among these, the first problem presents the most significant challenges due to the additional constraints imposed by the rigid power line on the robot’s motion, and thus serves as the primary focus of this paper in both simulation and real settings, illustrated in Fig. 2 (c) & (d). While our aim is not to provide an industry-grade system, it is noteworthy that no fully autonomous solutions currently exist for the power line problem. Additionally, we provide simulation videos for the agricultural exposure task to emphasise the generality of our approach. Formally, our contributions are as follows:
@@
    A novel context-agnostic occlusion heuristic and reward strategy generalizable to various exposure applications to clear deformables from a target region.
@@
Subsequent sections specifically focus on the power line clearance problem; however, the methodology is readily adaptable to others. We encourage readers to view the extensive supplementary videos to see our approach in action.
@@ II RELATED WORK
For deformable manipulation with high-dimensional object geometries, point clouds offer a rich sensor-agnostic scene mapping. Notable 1D implementations for tracking [16] and learning dense depth object descriptor features [17] leverage point representations. Point clouds have been used to learn policies with 2D and 3D deformables as well for cloth manipulation [8], liquid scooping [18], and dough cutting [7]. On the policy front, these deformable works use either Motion Planning with primitives [8], Task and Motion Planning [7], or Behavioural Cloning [18][17]. In contrast, model-free reinforcement learning works on deformables like ours are scarce, largely due to poor sample efficiency. With rigid body point clouds, model-free policies have been proposed [19][20] for interactive tasks, including sim-to-real transfer. Relevant to our setup, a key study with rigid bodies [21] found that RL with 3D point clouds outperforms its 2D image counterpart when object-object relationship encoding is crucial for the task.
@@
Post-scene capture, common deformable representation strategies include keypoints [16], graphs [22], and dense-descriptors [17]. Particularly relevant to our work, the deformable state can be represented as as a distributional embedding with a kernel mean operator [4]. They tackle a real-to-sim inference problem with an RKHS-net layer using a small number of keypoints ($\approx 10$) extracted from images. In contrast, we project entire point clouds ($\approx 512$) of multiple objects to RKHS and use the embeddings directly as RL observations in a sim-to-real context.
Our policy enables the arm to contort itself into shapes, such as hooks and props, leveraging the whole arm to push away the occlusion. While such formations are less prevalent, contact-aware reaching has been achieved with external full-body tactile sensors, for rigid clutter [9], or with internal sensors for tree branches [11]. However, we go beyond simple reach to sophisticated clearance tasks and implement a multi-modal policy integrating touch and vision feedback.
@@
Among deformables, interacting with plant structures is substantially harder owing to the non-uniformity in geometry, dynamics, and visual features. [23] proposes a method to gently grasp and move a leaf aside to estimate the pose of hidden fruits. On the other hand, [24] pushes away multiple leaves simultaneously with the end-effector to reveal the space behind them using large amounts of real-world data. Both works assume soft leafy resistance, which is typically addressed in farms with simpler tricks, such as blowing compressed air [25] to separate leaves and fruits. From a learning perspective, [26] and [27] employ vision-based RL for reaching vine pruning locations and for exposing fruits, respectively; however, both assume known plant topology while focusing on the end-effector motion. In comparison, our approach is designed to accommodate stiff resistance from multiple branches, uses the whole arm to maximise applied torques, and is trained in simulation without any digital twins, plant mesh similarity, or real-world data. Finally, none of the aforementioned works are designed for the hybrid rigid-deformable problem of power line clearance.
@@ III-A Deformable Simulation
Realistic simulation of the intricate tree-branch geometry is challenging; fortunately, this is well-studied in botanical literature under the L-system paradigm [28], predominantly for visualisation. An L-system borrows from formal grammar theory to model plant morphology by applying a recursive rule collection, starting from an axiom, mimicking the branching patterns. Recently, [11] has shown that this paradigm can be exported to 3D physics simulators, replete with realistic dynamics, by representing branch segments as rigid cylindrical links connected by revolute joints, amenable to robotic interaction. In this setup, branch deformations are modelled with a mass-spring-damper [29] actuated with proportional-derivative controllers. We borrow the recursion rules, growth attributes, and L-system parameters from [11]; however, we extend it with spherical joints to enable multi-axis deformation and enhance compliance for clearance, refer Fig. 2(d). We run thousands of such tree models in parallel on the distributed physics simulator Isaac Gym [30].
@@ III-B Policy Learning
We formulate the de-occlusion task as a discrete-time Markov Decision Process (MDP), where an agent learns a stochastic policy $\pi_{w}(a|o)$; parameterised by weights $w$. An MDP is defined by $\langle S,A,P_{a},R_{a},\beta\rangle$, where $S$ is states, $A$ is actions, $P_{a}$ and $R_{a}$ are the state transition model and the reward from action $a\in A$, and $\beta$ is a discount factor. The agent aims to maximise $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\beta^{t}R_{a}(s_{t},a_{t})\right]$, the expected discounted rewards over $T$ steps. Crucially, the agent only has access to noisy observations $o\in O$; for instance, point clouds from the camera or robot sensor measurements rather than the true state. Our algorithm choice is Proximal Policy Optimisation [31] from RL Games [32].
III-C Kernel Mean Embedding
In this section, we briefly examine distribution embedding and its spectral approximation before applying it to deformables; for a thorough explanation, see [33] and [34].
An RKHS is a Hilbert space of functions uniquely determined by a positive definite kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, that guarantees both an implicit feature mapping $\varphi:\mathcal{X}\to\mathcal{H}$ from the input space and the higher-dimensional target space $\mathcal{H}$. Akin to point mappings, a kernel mean operator lifts a probability distribution $\mathbb{P}$ to a single mean function in $\mathcal{H}$, however, without any loss of information for a characteristic kernel such as RBF: $k(x,x^{\prime})=\exp(-\frac{||x-x^{\prime}||^{2}}{2\gamma{}^{2}})$. We imagine $x$ to be points from the 3D space of point clouds. The lifted kernel mean is defined as:
[…]
While the inner product $\langle.,.\rangle_{\mathcal{H}}$ in eqn. (2) can indeed be computed with a finite-dimensional gram matrix, leveraging the reproducing property $\langle\varphi(x_{i}),\varphi(x_{j})\rangle_{\mathcal{H}}=k(x_{i},x_{j})$, but this is still problematic for large N. On the other hand, representing kernel functions through their spectral representations, particularly with Random Fourier Features (RFF) [34], significantly reduces the computational overhead. Through Bochner’s theorem, any kernel can be defined as a Fourier transform of a non-negative measure, subject to a stationarity constraint $k(x,x^{\prime})=k(x-x^{\prime})$.
We choose $R=16$ and use the RBF kernel, which results from the Fourier transform of a Gaussian $p(\omega)\sim\mathcal{N}(0,1)$.
III-D Distribution Embedding for Deformables
Given a segmented point cloud $P_{t}\in\mathbb{R}^{N\times 3}$ at time-step $t$, our spatial abstraction assumes it as samples from an underlying distribution, regardless of the deformability. The points in 3D coordinate space are elevated to the feature space $\mathcal{H}$ with an RFF approximated kernel mean, as described in Section III-C, to form the key constituent of the observation space $O$. We capture four independent point clouds of varying sizes: $P^{rob}$, $P^{clr}$, $P^{wbr}$, $P^{zbr}$, representing the robot, clearance region (i.e. a slice of the power line), the whole tree, and a zoomed-in version of the branches near the occlusion. The robot points are sampled from the URDF mesh based on link positions at each time step. Points representing the cylindrical clearance region are sampled from the cylinder surface, given its fixed position and orientation. In contrast, camera sensors provide the deformable and time-varying points of the occluding branches.
There are several advantages to such a representation. Firstly, although the explicitly mapped feature locations are unknown, the agent can exploit the relative similarity of embedded points in RKHS, for instance, to minimise the foliage-power line entanglement or maximise the gripper-branch proximity, resembling the ‘kernel trick’ common in machine learning literature. Secondly, a distributional interpretation is invariant to point permutations and robust to the high noise content characteristic of streaming point clouds. Third, unlike conventional point cloud feature extractors like PointNet++ [35], kernel embeddings are resilient to variations in the size and structure of input point clouds, amplified by the arm, branches and the power line frequently occluding each other. Finally, and most crucially, this approach is significantly faster, reducing training time and enabling real-time inference, as our experiments demonstrate. This speed advantage stems from formulation in Section III-C, where the feature map $\hat{\varphi}(x)$ can be derived with a single matrix product of the input points $x$ and the pre-computed Fourier coefficients ${\omega}_{r}^{\top}$ of complexity $\mathcal{O}(N\times R)$, followed by a sin/cos transformation, and an additional row-average to obtain the embedding $\overline{\mu_{\mathbb{P}}}$, a highly efficient pipeline scalable to thousands of environments, amenable to GPU parallelisation.
III-E Occlusion Heuristic
@@
A key consideration in the de-occlusion problem is determining an effective yet simple measure to capture the temporal progression in occlusion level from manipulation. Common metrics include visible surface ratio, pixel counts [36] (e.g., red fruit pixels in a green foliage background), or labelled occlusion levels. However, these metrics depend either on prior knowledge of the target geometry or partial observability at start, and they may not be smooth enough for an RL agent. In contrast, our occlusion heuristic $h_{t}\in[0,1]$ is independent of target characteristics or even its presence.
@@
$h(P^{(1)},P^{(2)}):=\frac{1}{k}\sum_{d\in\mathcal{D}_{k}}\mathbb{I}(d<d_{\text{th}}).$
Intuitively, it is the fraction of point pairs (from different groups) that breach a proximity threshold to the total number of pairs in the neighbourhood, where both the threshold $d_{th}$ and the neighbourhood size $k$ are tunable. In general, the larger the number of branch segments covering the clearance region, the higher the $h_{t}$. We set $k=200$ and the threshold to be the clearance region cylinder diameter $d_{th}=10$cm. We emphasise that the clearance region has no shape constraints due to its point representation; for instance, it can be defined to de-occlude an expected curved end-effector trajectory of a second manipulator in a dual-arm setting.
III-F Observation & Reward Space

### v2

@@ Abstract
Website: https://sites.google.com/view/dcmwap/
@@ I INTRODUCTION
Learning to manipulate deformable objects poses a formidable challenge in robotics [yin2021modeling]; however, operating with clusters of deformables, such as cable bundles, multi-branch tree canopies, and cloth piles, is significantly harder. Although imprecise physics knowledge [arriola2020modeling] and sensor uncertainties are contributing factors, the principal challenge lies in inferring a compact, low-dimensional state abstraction and the temporal state evolution, based on noisy perception from a high-dimensional ground truth [antonova2022bayesian]. Furthermore, with deformable clusters, tiny differences in contact locations, material properties, and initial configuration among the individual cluster constituents (e.g. between thicker and thinner plant stems) give rise to non-linear resistance forces and divergent motion trajectories upon interaction. The resulting chaotic dynamics render model-based learning infeasible, while model-free solutions have high sample complexity, and therefore, the compactness and generation efficiency of the state representation gather significance [maillard2011selecting].
@@
Novel deformable works manipulating cables [grannen2020untangling], dough [lin2022planning], and clothes [deng2024general] are all planning approaches; furthermore, they exclusively focus on end-effector-based control. In a cluster context, such grasp-based solutions tailored to individual deformable members become prohibitively long. On the contrary, a multi-link manipulation strategy substantially escalates the contact complexity, necessitating additional perception modalities apart from vision. However, typical auxiliary sensors for whole-arm contact detection are either expensive [jain2013reaching] or too slow [huang20243d] for low-latency applications.
Addressing these challenges, we propose a model-free, non-prehensile, multi-modal, reinforcement learning (RL) approach operating on the cluster as a whole with full-body contact, but without external tactile sensing. We use a distributional representation [antonova2022bayesian] of the cluster state generated by embedding point cloud into a Reproducing Kernel Hilbert Space (RKHS), without key-point intermediaries. Compared to traditional point cloud feature extractors, a distribution embedding is notably lightweight, enabling faster training and real-time inference. It eliminates the need for points to track identical cluster locations, is invariant to permutations, and can handle variable input sizes due to members moving out of camera view between frames. Further, the observations include a whole-arm contact classifier [jacobgentle] output from proprioceptive input to flag contacts. Finally, we address the high sample requirements of on-policy RL by leveraging the parallelism of physics simulators during training and transferring it zero-shot to the real-world.
Refer to caption Refer to caption
(a) (b)
Figure 2: (a) Manual pruning of overhanging branches impeding power lines, from [siebert2014survey] (b) Our simulation setup with an L-system structure for the power line clearance scenario.
This work focuses on a specific subclass of problems, aiming to de-occlude local regions of the deformable cluster; for instance, to clear a cylinder-shaped region among tree branches, by manipulating a few of them together. We implement two illustrative applications of such a skill abstraction:
@@
    Clearing Power Lines: Hazardous tree branches encroaching on overhead power infrastructure is a leading cause of forest fires and outages [lowe2022identification][guggenmoos2003effects]. Pruning foliage near overhead lines (Fig. 2a). is dangerous for humans due to risks like falls, shocks, and tool accidents [siebert2014survey]; therefore, robotic assistive systems that can clear the branches away from the lines before pruning can enhance safety while reducing outages.
@@
    Agricultural Exposure: Modern harvesting systems suffer from lower success rates with partially (50%-75%) and fully occluded (5%) fruits [zhou2022intelligent]. Manipulating plant stems for exposing fruits or a diseased region to an external inspection camera requires a line-of-sight restoration through multi-branch clearance.
Among these, the first problem presents the greatest challenge due to additional constraints imposed by the rigid power line on robot’s motion, and thus serves as the primary focus of this paper in both simulation and real settings. While our aim is not to provide an industry-grade system, it is notable that no autonomous solutions currently exist for the power line problem. Additionally, we provide simulation videos for the agricultural exposure task to emphasise the generality of our method. Specifically, our contributions are:
@@
    A novel context-agnostic occlusion heuristic and reward strategy generalisable to various exposure applications to clear deformables from a target region.
@@
Our first two contributions are technical and apply to deformable manipulation in general, while the third and fourth are system-level, focusing on de-occlusion, but deployable beyond trees. Subsequent sections specifically focus on the power line clearance problem; however, the methodology is readily adaptable to others. We encourage readers to view the supplementary videos to see our approach in action.
@@ II RELATED WORK
For deformable manipulation with high-dimensional geometries, point clouds offer a rich sensor-agnostic scene mapping. Notable 1D implementations for tracking [tang2022track] and learning dense depth object descriptor features [sundaresan2020learning] leverage point representations. Point clouds have been used to learn policies with 2D and 3D deformables as well for cloth manipulation [deng2024general], liquid scooping [seita2023toolflownet], and dough cutting [lin2022planning]. On the policy front, these deformable works use either Motion Planning with primitives [deng2024general], Task and Motion Planning [lin2022planning], or Behavioural Cloning [sundaresan2020learning]. In contrast, model-free RL works on deformables like ours are scarce, largely due to poor sample efficiency. With rigid body point clouds, model-free policies have been proposed [liu2022frame] for interactive tasks, including sim-to-real transfer.
@@
Post-scene capture, deformable representation strategies include keypoints [tang2022track], graphs [ma2021learning], and dense-descriptors [sundaresan2020learning]. Relevant to our work, the deformable state can be represented as as a distributional embedding with a kernel mean operator [antonova2022bayesian]. They tackle a real-to-sim inference problem with an RKHS-net layer using a small number of keypoints ($\approx 10$) extracted from images. In contrast, we project entire point clouds ($\approx 512$) of multiple objects to RKHS and use the embeddings directly as RL observations in a sim-to-real context. Our policy enables the arm to contort itself into hooks and props, leveraging the whole arm for clearance. While such formations are less prevalent, contact-aware reaching has been achieved with external full-body tactile sensors, for rigid clutter [jain2013reaching], or with internal sensors for tree branches [jacobgentle]. However, we go beyond simple reach to sophisticated clearance tasks and implement a multi-modal policy integrating touch and vision feedback.
@@
Among deformables, plant interaction is substantially harder owing to the non-uniformity in geometry, dynamics, and visual features. While [yao2024safe] gently moves a leaf aside to estimate hidden fruit pose, [zhang2023push] clears multiple leaves simultaneously with the end-effector using large quantities of real-world data. Both works assume soft leafy resistance, which is typically addressed in farms with simpler tricks, such as blowing compressed air [yamamoto2014development] to separate leaves and fruits. From a learning perspective, [yandun2021reaching] and [subedi2025find] employ vision-based RL for reaching vine pruning locations and for exposing fruits, respectively; however, both assume known plant topology while focusing on the end-effector. In comparison, our approach accommodates stiff resistance from multiple branches, uses the whole arm to maximise applied torques, and is trained in simulation without digital twins, plant mesh similarity, or real-world data. Finally, none of the aforementioned works are designed for the hybrid rigid-deformable problem of power line clearance.
@@ III APPROACH
A brief overview of our approach (Fig. 1) is as follows. First, we generate parallel simulations of deformable tree-branches (Sec. III-A) to train an RL policy (Sec. III-B). To guide policy learning, we compute an occlusion metric from the scene, which serves as a reward signal (Sec. III-C). The policy is trained in a physics simulator with domain randomisation applied to the geometry and dynamics of trees (Sec. III-D). The learning framework leverages a kernel mean operator (Sec. III-E) to encode the segmented point clouds as RL features (Sec. III-F). In addition, the features include robot metrics and the output of a touch classifier. For sim-to-real transfer (Sec. IV-B), we use several vision-based segmentation and masking methods to process camera images, generate object masks and extract the vision features necessary for real-world policy deployment (Sec. IV-C).
@@ III-A Deformable Simulation
Realistic simulation of the intricate tree-branch geometry is challenging; fortunately, this is well-studied in literature under the L-system paradigm [prusinkiewicz2012algorithmic], predominantly for visualisation. An L-system borrows from formal grammar theory to model plant morphology by applying a recursive rule collection, starting from an axiom, mimicking the branching patterns. Recently, [jacobgentle] has shown that this paradigm can be exported to 3D physics simulators, by representing branch segments as rigid cylindrical links connected by revolute joints, amenable to robotic interaction. In this setup, branch deformations are modelled with a mass-spring-damper [jacob2024learning] actuated with proportional-derivative controllers. We borrow the recursion rules, growth attributes, and L-system parameters from [jacobgentle]; however, we extend it with spherical joints to enable multi-axis deformation and enhance compliance, refer Fig. 2(b). We run thousands of such tree models in parallel on the distributed physics simulator Isaac Gym [makoviychuk2021isaac].
@@ III-B Policy Learning
We formulate the de-occlusion task as a discrete-time Markov Decision Process (MDP), where an agent learns a stochastic policy $\pi_{w}(a|o)$; parameterised by weights $w$. An MDP is defined by $\langle S,A,P_{a},R_{a},\beta\rangle$, where $S$ is states, $A$ is actions, $P_{a}$ and $R_{a}$ are the state transition model and the reward from action $a\in A$, and $\beta$ is a discount factor. The agent aims to maximise $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\beta^{t}R_{a}(s_{t},a_{t})\right]$, the expected discounted rewards over $T$ steps. Crucially, the agent only has access to noisy observations $o\in O$; for instance, point clouds from the camera or robot sensor measurements rather than the true state. Our RL algorithm choice is Proximal Policy Optimisation [schulman2017proximal].
III-C Occlusion Heuristic as Reward Signal
@@
A key consideration for de-occlusion is determining an effective yet simple measure to capture the temporal progression in occlusion level from manipulation, for instance, as a reward signal to guide the policy learning. Common metrics include visible surface ratio, pixel counts [bai2025retrieval] or labelled occlusion levels. However, these metrics depend either on prior knowledge of the target geometry or partial observability at start, and may not be smooth enough for an RL agent. In contrast, our occlusion heuristic $h_{t}\in[0,1]$ is independent of target characteristics or its presence.
@@
$h(P^{(1)},P^{(2)}):=\frac{1}{k}\sum_{d\in\mathcal{D}_{k}}\mathbb{I}(d<d_{\text{th}}).$ (1)
Intuitively, it is the fraction of point pairs (from different groups) that breach a proximity threshold to the total number of pairs in the neighbourhood, where both the threshold $d_{th}$ and the neighbourhood size $k$ are tunable. Larger the number of branch segments covering the clearance region, higher the $h_{t}$. The threshold $d_{th}$ (set to 10cm) is the safe distance for an auxiliary robot to prune the cleared branches. The $k$ (set to 200) balances the branch-line interaction fidelity against computational cost; a lower $k$ risks skipping critical contact points, while a higher $k$ significantly increases processing time. We emphasise the clearance region has no shape constraints due to its point representation; for instance, it can be defined to de-occlude an expected curved end-effector trajectory of a second manipulator in a dual-arm setting.
III-D Domain Randomisation for RL training
This work aims to generalise a learned policy to deformables with unseen geometry, undetermined dynamics, and devoid of chromatic affinity, by not relying on digital twins [subedi2025find], parameter inference [jacob2024learning][antonova2022bayesian], or RGB images during training. Instead, we perturb simulation parameters, for example, L-system traits such as branch divergence angle and elongation rate, and dynamics parameters such as spring stiffness and damping. Further, we subsample and permute all point clouds while adding noise at each time-step.
III-E Kernel Mean Embedding for Observation Encoding
[…]
$\varphi(\mathbb{P})=\mu_{\mathbb{P}}:=\mathbb{E}_{x\sim\mathbb{P}}[k(\cdot,x)]=\int k(\cdot,x)\,d\mathbb{P}$ (2)
While the inner product $\langle.,.\rangle_{\mathcal{H}}$ in eqn. (3) can indeed be computed with a finite-dimensional gram matrix, leveraging the reproducing property $\langle\varphi(x_{i}),\varphi(x_{j})\rangle_{\mathcal{H}}=k(x_{i},x_{j})$, but this is still problematic for large N. On the other hand, representing kernel functions through their spectral representations, particularly with Random Fourier Features (RFF) [rahimi2007random], significantly reduces the computational overhead. Through Bochner’s theorem, any kernel can be defined as a Fourier transform of a non-negative measure, subject to a stationarity constraint $k(x,x^{\prime})=k(x-x^{\prime})$.
A higher $R$ indicate more informative features, but an excessive value explodes the observation space, demanding more samples during training. As a trade-off, we choose $R=16$ and use the RBF kernel, which results from the Fourier transform of a Gaussian $p(\omega)\sim\mathcal{N}(0,1)$.
III-F Distribution Embedding for Deformable Clearance
Given a point cloud $P_{t}\in\mathbb{R}^{N\times 3}$ at time-step $t$, we assume it as samples from an underlying distribution. These 3D points are elevated to RKHS with an RFF approximated kernel mean to form the key constituent of RL observations. We capture four independent point clouds of varying sizes: $P^{rob}$, $P^{clr}$, $P^{wbr}$, $P^{zbr}$, representing the robot, clearance region (i.e. a slice of the power line), the whole tree, and a zoomed-in version of the branches near the occlusion. While the robot points are sampled from URDF mesh based on link positions at each step, the cylindrical clearance region points are sampled from its surface, given a fixed pose. In contrast, camera sensors provide the deformable and time-varying points of the occluding branches.
There are several advantages to such a representation. Firstly, although the explicitly mapped feature locations are unknown, the agent can exploit the relative similarity of embedded points in RKHS, for instance, to minimise the foliage-power line entanglement or maximise the gripper-branch proximity, resembling the ‘kernel trick’ common in machine learning. Secondly, a distributional interpretation is invariant to point permutations and robust to the high noise content characteristic of streaming point clouds. Third, unlike conventional point cloud feature extractors like PointNet++ [qi2017pointnet++], kernel embeddings are resilient to variations in the size and structure of input point clouds, amplified by the arm, branches and the power line frequently occluding each other. Finally, and most crucially, this approach is significantly faster, reducing training time and enabling real-time inference, as our experiments demonstrate. This speed advantage stems from formulation in Section III-E, where the feature map $\hat{\varphi}(x)$ can be derived with a single matrix product of the input points $x$ and the pre-computed Fourier coefficients ${\omega}_{r}^{\top}$ of complexity $\mathcal{O}(N\times R)$, followed by a sin/cos transformation, and an additional row-average to obtain the embedding $\overline{\mu_{\mathbb{P}}}$, a highly efficient pipeline scalable to thousands of environments, amenable to GPU parallelisation.
III-G Observation & Reward Space